OpenStudy (anonymous):

Simple Arithmetic Problem: Using any Maths Symbols or Signs prove it... 0 0 0 = 6 1 1 1 = 6 2 2 2 = 6 3 3 3 = 6 4 4 4 = 6 5 5 5 = 6 6 6 6 = 6 7 7 7 = 6 8 8 8 = 6 9 9 9 = 6 You can Insert Everything but Numbers, And you can't change everything.

OpenStudy (rsadhvika):

\(\ne\)

OpenStudy (rsadhvika):

can i use that sign

OpenStudy (anonymous):

no. just math operations

OpenStudy (goldphenoix):

6-6+6 = 6 :D

OpenStudy (anonymous):

that's one

OpenStudy (goldphenoix):

2+2+2=6. :D

OpenStudy (anonymous):

that's two

OpenStudy (goldphenoix):

3*3-3=6. :D I got 3 points so far!

OpenStudy (shubhamsrg):

( 0! + 0! + 0!)! = 6 ( 1 + 1 +1) ! =6

OpenStudy (anonymous):

good good keep it coming

OpenStudy (shubhamsrg):

signum function and successor function will take us anywhere! :D but we have to use only operators..hmm

OpenStudy (anonymous):

4 down

OpenStudy (goldphenoix):

Square root of 4 + Square root of 4+ Square root of 4 = 6 :D

ganeshie8 (ganeshie8):

(3!- 3+ 3) = 6

OpenStudy (anonymous):

5 down

OpenStudy (anonymous):

that seems legit

OpenStudy (anonymous):

6

OpenStudy (shubhamsrg):

6*6/6 = 6

OpenStudy (shubhamsrg):

7- (7/7)

OpenStudy (goldphenoix):

5+(5/5) = 6 :D

OpenStudy (anonymous):

i think 8 and 9 are the last

OpenStudy (skullpatrol):

Where in math will this help?

OpenStudy (anonymous):

@skullpatrol more than you know. many top users disappeared because of the disappearance of these things. and the disappearance of the top users became the disappearance of the good helpers here. but i wouldn't want to delve on those subjective matters

ganeshie8 (ganeshie8):

\(\log_2 4 + \log_2 4 + \log_2 4\) = 6

OpenStudy (anonymous):

lol wow. logs got in

OpenStudy (anonymous):

but i don't think that's legit

OpenStudy (anonymous):

you put in numbers

OpenStudy (shubhamsrg):

also in sqrt4 + sqrt4 + sqrt4 , wont you consider "2" being used ?

OpenStudy (shubhamsrg):

in 4^(1/2)

OpenStudy (anonymous):

sqrt is a math symbol...but if you want to reject that then that leaves 4 8 and 9

OpenStudy (goldphenoix):

Cube root of 8 + Cube root of 8+ Cube Root of 8 = 6 :D I'm on the role!

OpenStudy (shubhamsrg):

well yes cube root can also be taken into account, and then any root should be taken into account ?

OpenStudy (anonymous):

shubhamsburg disqualified roots though

OpenStudy (anonymous):

that was an extra bu

OpenStudy (goldphenoix):

Aww man.

ganeshie8 (ganeshie8):

oh roots we cant abuse

OpenStudy (unklerhaukus):

sometime you can write binary log (ie base 2) \( \log_2(x) =\text{lb}(x)\)

OpenStudy (anonymous):

lol..

OpenStudy (shubhamsrg):

i'd prefer ( 4- (4/4))!

OpenStudy (anonymous):

8 and 9 then

OpenStudy (skullpatrol):

I disqualify this question as having no use in math.

OpenStudy (goldphenoix):

It's like saying that algebra has no use in math. :|

OpenStudy (anonymous):

8 and 9 seem difficult without roots

OpenStudy (shubhamsrg):

we can use operators hmm how about ++ and -- ? :P

OpenStudy (anonymous):

isn't that a logical operator?

OpenStudy (shubhamsrg):

not a genuine operator we use in theoretical maths.hmm we use that in computer programs

OpenStudy (shubhamsrg):

nvm

OpenStudy (shubhamsrg):

i dont see anything for 8 or 9

OpenStudy (anonymous):

because you disqualified the roots :p lol

OpenStudy (goldphenoix):

Use the power of square root! :|

OpenStudy (anonymous):

since you answered 8 with roots how would you answer 9 @GoldPhenoix

OpenStudy (anonymous):

here is my rather late and lame contribution \[\sqrt{9}!+9-9\]

OpenStudy (anonymous):

this has to be the best question i have seen all week

OpenStudy (goldphenoix):

Not bad, satellite.

OpenStudy (anonymous):

it's nice to have some math riddles sometimes

OpenStudy (goldphenoix):

Yes, totally agree.

OpenStudy (skullpatrol):

Ok, I'll agree it is "recreational" arithmetic :)

OpenStudy (unklerhaukus):

10+10+10=6 (binary)

OpenStudy (shubhamsrg):

right! :D

ganeshie8 (ganeshie8):

11 + 11 + 11 = 6 (unary)

OpenStudy (unklerhaukus):

nice

OpenStudy (shubhamsrg):

reminds me of a quote, "there are only 10 types of people in the world, one who knows binary, other who doesn't! " :P

OpenStudy (anonymous):

lol

OpenStudy (skullpatrol):

Thanks for the question :D